# AI sustains higher strategic tension than humans in chess

### v2

@@ Abstract
The aphorism that one may have won the battle but lost the war is encapsulated in the notion of a “Pyrrhic victory.” Costly short-term gains must be balanced against the longer-term uncertainties, opportunities, or challenges that may emerge in competitive environments. This presents a fundamental challenge in strategic decision-making for both biological and artificial agents across diverse competitive and cooperative landscapes [42, 40, 3], of which war [weinsteinRebellionPolitics2007, bohorquezCommonEcology2009] and conflicts [smithLogicAnimal1973, dedeoInductiveGame2010a, brushConflictsInterest2018] are poignant examples. In the more general case, organisms and individuals learn such as during development (e.g., neural pruning [woudeNoGains2019]), foraging (e.g., depending on climate [piteraDailyForaging2018]), and social change (e.g., reorganization of social hierarchy [28]) or, over evolutionary time scales, “adapt” to changing ecological niches [24]. In the context of an individual’s lifetime, these can be characterized as strategic choices that involve an implicit trade-off between overfitting the local environment and maintaining the flexibility to adapt to unforeseen environmental conditions [27], in an echo of the speed-accuracy trade-off [33]. In many real-world settings, a complete measurement of the strategic environment, costs, and outcomes is not available. Chess, by contrast, is a game of perfect information, providing a well-studied and bounded setting.
@@
As a result, chess has served as a model system for investigating strategy. It has well-defined rules, perfect information, complexity, and rich history as a testbed for both human expertise and artificial intelligence (AI) [36, 39, 11, 32]. Chess is a competitive game between two players, white and black, who take turns making moves with a single piece on a board with 8x8 squares as shown in Fig. 1. Both players start with 16 pieces, and the objective of the game is to capture the enemy’s king. A key feature that makes the game interesting is the combinatorial explosion in potential moves, which prohibits exhaustive calculation. As a result, human players must make heuristic decisions about the possible long-term value of moves [39]. The nature of the trade-off is shaped by cognitive limitations, psychological pressures, and time constraints, often leading to biased or suboptimal decisions [21, 25, 20]. In contrast, modern AI exhibits superhuman performance through two main strategies. One class involves deep reinforcement learning systems, such as AlphaZero and Leela Chess Zero, which develop strategies through self-play using deep convolutional neural networks. Alongside these, engines like Stockfish combine massive brute-force search trees with NNUE (Efficiently Updatable Neural Networks). While Stockfish originally relied on handcrafted evaluation functions, modern versions utilize machine learning to train their evaluation layers, blending classical search efficiency with neural network-based intuition. These differences suggest that the nature of the trade-offs that AI engines make is distinct from those made by human players. Still, there is little characterization of the properties that distinguish gameplay.
@@ I Defining Tension in Chess
In constructing the network, we disregard which side is to move, focusing instead on the full set of interactions at each ply to capture the full state of the board. Attack and defense links can have weight 1 or 2, depending on whether the corresponding conditions are satisfied by one or both of the pieces involved. In contrast, control links always have weight 1. By construction, the network has at most 64 connected nodes, each corresponding to either a chess piece or a vacant square. For our work, we do not distinguish between link types  [^1] Although we also experimented with alternative networks that distinguish between link types and assign values according to the pieces involved, following commonly accepted conventions, these variations yielded qualitatively similar results in our case studies. We therefore opted for the simpler representation to maintain interpretability and focus on the main structural features of the positions.. The resulting network captures the interplay among pieces and squares that creates and sustains tension on the chessboard.
Figure 1: One example of tension network. (a) Chessboard at the ply 44 during a game between the two grandmasters, Magnus Carlsen (white) and Alexey Sarana (black), in the Titled Tuesday tournament, Round 6, held on chess.com on November 26, 2024. (b) The network represents the interactions on the chessboard. The chess pieces are depicted as dark and light circles, depending on whether they are black or white. Attack links are displayed in orange when originating from white pieces and in red when originating from black. Defense links are turquoise when white is defending and blue otherwise. Control links are light blue (originating white) and light green (originating black). Small, brownish nodes represent vacant squares under control. The chosen ply represents the highest tension reached during the game according to the maximum eigenvalue of the adjacency matrix.
@@
We illustrate the typical evolution of the strategic tension across multiple games in Fig. 2. We can distinguish four main phases. The first phase, corresponding roughly to the opening stage up to around ply 30, features a buildup where pieces approach each other, and contested squares come under threat, leading to an abrupt increase in tension. The second phase, spanning approximately from ply 30 to ply 60, represents peak complexity, where the board remains highly interconnected, and the tension stays elevated. The third phase exhibits a gradual simplification of the tension network, as piece exchange and loss steadily lower the overall tension. Finally, the games reach a relatively stable value of tension, but at this point available statistics become increasingly sparse, especially for games played between humans that typically end earlier (Fig. 2b and 2c). Thus, the dynamics of tension capture the overall contours of the game, including the opening buildup, the combinatorially complex midgame, and the relaxation leading up to the endgame as discussed in the theory of chess play [13].
Despite the overall similarities in the typical trajectory of strategic tension between human and AI play, there are notable differences. Maximum tension is higher in human-played games than in engine-played games. Up to approximately ply 40, human players typically follow established opening repertoires from theoretical databases, which often lead to well-studied positions with numerous direct piece interactions, including attacks, defenses, and immediate threats [15, 6]. These theory-driven positions generate the highest points of tension as players transition to original play in the midgame. Once players leave the realm of memorized theory and must rely on their own calculation and evaluation, they tend to simplify positions. Human players reduce tension more rapidly and, as we discuss below, resolve piece conflicts sooner, often favoring attack-oriented configurations. In contrast, state-of-the-art engines tend to maintain more distributed, positional advantages that unfold gradually over multiple moves, resulting in a slower decay of tension.
@@ II Result 2: Strategic tension reflects, does not predict game result
To relate the structural complexity of the interaction network to game outcomes, we compare the evolution of strategic tension with the algorithmic evaluation provided by Stockfish (see Fig. 8 in the Supplementary Materials). Inspecting the evaluation score, we identify bifurcation points in game trajectories, typically around ply 30 (move 15), where decisive and drawn games begin to diverge for human players during the opening-to-midgame transition. A similar behavior is observed in AI games, where a local maximum or inflection point in the evaluation profile precedes the divergence of outcomes. Notably, in all games, this divergence occurs before strategic tension reaches its absolute peak. This ordering suggests that the game’s outcome is largely determined during the ascent toward peak tension, before the subsequent reduction in tension differentiates decisive games from draws. In chess terms, foundational advantages—such as positional pressure, space, or superior piece coordination—are established while the board is still increasing in complexity.
@@
Attack-defense balance: When distinguishing between attack and defense links, we find that the surplus of attack links over defense links, or “attack-defense balance” as shown in Fig. 4, moves inversely with the tension profile. It first drops in the early phase more defense-oriented phase of the game. As the game progresses, attack links gradually outnumber defense links. The curve is shifted towards more positive values in human games compared to AI games.
@@
Figure 5: Tension loads for different Elo ratings and Stockfish depths. (a) We represent tension loads (sum of tension values up to ply 150) for rapid games (turquoise dots), blitz games (green dots), classical games played by top human players (blue star), and games between Stockfish and Leela from official TCEC tournaments (red star), all associated with estimated Elo ratings. Each dot/star represents the average across 1,200 games. The blue and red stars were derived from the data analyzed in Fig. 2. For reference, blitz games are played with very short time controls (typically 3–5 minutes per player), rapid games with intermediate time controls (around 10–25 minutes per player), and classical games with long time controls (often 60 minutes or more per player) allowing deeper strategic play. (b) We represent tension loads from games between versions of Stockfish at varying depth levels (orange dots), computed from a smaller sample of 120 games per dot. The two panels also show the $R^{2}$ values of the corresponding linear regressions.
While engines can maintain higher tension for a longer duration compared to humans, does this difference depend on skill level? We consider the cumulative tension up to ply 150 across different Elo ratings for human players in Fig. 5a  [^3] In principle, Stockfish has a search depth parameter that serves as a proxy for skill level. Leela has a time limit parameter during which it is permitted to assess the board state and make a decision. At a few dozen search depths, the tension load is limited to about 450, similarly for Leela at a few seconds. Since the games that are shown go far beyond these limited parameter values with supercomputers, this suggests that algorithmic complexity increases the ability to handle tension.. Tension changes substantially with Elo, showing an approximately linear dependence in human games. Moreover, across all Elo values, rapid games consistently exhibit slightly higher tension loads than blitz games. This is consistent with the idea that longer time controls allow human players to create and manage more complex positions [5]. The effect is even more pronounced in classical games, where players have significantly more time to make decisions, resulting in higher tension and reflecting the chess community’s concern that time-constrained players experience larger performance drops on complex decisions and therefore tend to choose safer, less strategically risky continuations [5, 7]. In contrast, elite classical play is often characterized by keeping simplification at bay and maintaining a complex struggle, where grandmasters sustain tense, fluid positions that test the opponent’s stamina and depth of calculation over many hours [22, 30, 12].
@@ III Discussion
AC led the definition of the tension metric, with contributions from the other authors, and carried out the code development, numerical analyses, and figure generation. EDL performed several simulations on a computing cluster using Stockfish at high depths. VDPS provided the original inspiration for the project, supervised its development at all stages, and independently cross-checked the results. All authors participated in the discussions and contributed to writing the manuscript.
@@ Acknowledgements
We thank V. Latora for valuable discussions, particularly regarding the inclusion of Fig. 4, and M. Barthelemy for useful discussions and for sharing his code, which we used as a baseline for our analysis. This work was partly funded by the City of Vienna, Municipal Department 7, and the Federal Ministry of the Republic of Austria for Climate Action, Environment, Energy, Mobility, Innovation, and Technology as part of the project GZ2021-0.664.668. It was also partially funded by the Austrian Science Fund (FWF) 10.55776/ESP127.

### v3

@@ Abstract
The aphorism that one may have won the battle but lost the war is encapsulated in the notion of a “Pyrrhic victory.” Costly short-term gains must be balanced against the longer-term uncertainties, opportunities, or challenges that may emerge in competitive environments. This presents a fundamental challenge in strategic decision-making for both biological and artificial agents across diverse competitive and cooperative landscapes [42, 40, 3], of which war [weinsteinRebellionPolitics2007, bohorquezCommonEcology2009] and conflicts [smithLogicAnimal1973, dedeoInductiveGame2010a, brushConflictsInterest2018] are poignant examples. In the more general case, organisms and individuals learn such as during development (e.g., neural pruning [woudeNoGains2019]), foraging (e.g., depending on climate [piteraDailyForaging2018]), and social change (e.g., reorganization of social hierarchy [28]) or, over evolutionary time scales, “adapt” to changing ecological niches [24]. In the context of an individual’s lifetime, these can be characterized as strategic choices that involve an implicit trade-off between overfitting the local environment and maintaining the flexibility to adapt to unforeseen environmental conditions [27], in an echo of the speed-accuracy trade-off [33]. In many real-world settings, complete measurements of the strategic environment, costs, and outcomes are not available. Chess, by contrast, is a game of perfect information, offering a well-studied, bounded setting.
@@
As a result, chess has served as a model system for investigating strategy. It has well-defined rules, complexity, and rich history as a testbed for both human expertise and artificial intelligence (AI) [36, 39, 11, 32]. Chess is a competitive game between two players, white and black, who take turns making moves with a single piece on a board with 8x8 squares as shown in Fig. 1. Both players start with 16 pieces, and the objective of the game is to capture the enemy’s king. A key feature that makes the game interesting is the combinatorial explosion in potential moves, which prohibits exhaustive calculation. As a result, human players must make heuristic decisions about the possible long-term value of moves [39]. The nature of the trade-off is shaped by cognitive limitations, psychological pressures, and time constraints, often leading to biased or suboptimal decisions [21, 25, 20]. In contrast, modern AI exhibits superhuman performance through two main strategies. One class involves deep reinforcement learning systems, such as AlphaZero and Leela Chess Zero, which develop strategies through self-play using deep convolutional neural networks. Alongside these, engines like Stockfish combine massive brute-force search trees with NNUE (Efficiently Updatable Neural Networks). While Stockfish originally relied on handcrafted evaluation functions, modern versions utilize machine learning to train their evaluation layers, blending classical search efficiency with neural network-based intuition. These differences suggest that the nature of the trade-offs that AI engines make is distinct from those made by human players. Still, there is little characterization of the properties that distinguish gameplay.
@@ I Defining Tension in Chess
In constructing the network, we disregard which side moves, focusing instead on the full set of interactions at each ply to capture the board’s full state. Attack and defense links can have weight 1 or 2, depending on whether the corresponding conditions are satisfied by one or both of the pieces involved. In contrast, control links always have weight 1. By construction, the network has at most 64 connected nodes, each corresponding to either a chess piece or a vacant square. For our work, we do not distinguish between link types  [^1] Although we also experimented with alternative networks that distinguish between link types and assign values according to the pieces involved, following commonly accepted conventions, these variations yielded qualitatively similar results in our case studies. We therefore opted for the simpler representation to maintain interpretability and focus on the main structural features of the positions.. The resulting network captures the interplay among pieces and squares that creates and sustains tension on the chessboard.
Figure 1: One example of tension network. (a) Chessboard at the ply 44 during a game between the two grandmasters, Magnus Carlsen (white) and Alexey Sarana (black), in the Titled Tuesday tournament, Round 6, held on chess.com on November 26, 2024. (b) The network represents the interactions on the chessboard. The chess pieces are depicted as dark or light circles, depending on their color. The attack links are shown in red when originating from black and in orange when originating from white. Defense links are turquoise when white is defending and blue otherwise. Control links are light blue (originating white) and light green (originating black). Small, brownish nodes represent vacant squares under control. The chosen ply represents the highest tension reached during the game according to the maximum eigenvalue of the adjacency matrix.
@@
We illustrate the typical evolution of the strategic tension across multiple games in Fig. 2. We can distinguish four main phases. The first phase, roughly corresponding to the opening stage up to around ply 30, features a buildup in which pieces approach each other and contested squares come under threat, leading to an abrupt increase in tension. The second phase, spanning approximately from ply 30 to ply 60, represents peak complexity, where the board remains highly interconnected, and the tension stays elevated. The third phase exhibits a gradual simplification of the tension network, as piece exchange and loss steadily lower the overall tension. Finally, the games reach a relatively stable value of tension, but at this point available statistics become increasingly sparse, especially for games played between humans that typically end earlier (Fig. 2b and 2c). Thus, the dynamics of tension capture the overall contours of the game, including the opening buildup, the combinatorially complex midgame, and the relaxation leading up to the endgame as discussed in the theory of chess play [13].
Despite the overall similarities in the typical trajectory of strategic tension between human and AI play, there are notable differences. Maximum tension is higher in human-played games than in engine-played games. Up to approximately ply 40, human players typically follow established opening repertoires from theoretical databases, which often lead to well-studied positions with numerous direct piece interactions, including attacks, defenses, and immediate threats [15, 6]. These theory-driven positions generate the highest points of tension as players transition to original play in the midgame. Once players leave the realm of memorized theory and must rely on their own calculation and evaluation, they tend to simplify positions. Human players reduce tension more rapidly and, as we discuss below, resolve piece conflicts sooner, often favoring attack-oriented configurations. In contrast, state-of-the-art engines tend to maintain more distributed positional advantages that unfold gradually over multiple moves, leading to slower tension decay.
@@ II Result 2: Strategic tension reflects, does not predict game result
To relate the structural complexity of the interaction network to game outcomes, we compare the evolution of strategic tension with the algorithmic evaluation provided by Stockfish (see Fig. 8 in the Supplementary Materials). Inspecting the evaluation score, we identify bifurcation points in game trajectories, typically around ply 30 (move 15), where decisive and drawn games begin to diverge for human players during the opening-to-midgame transition. A similar behavior is observed in AI games, where a local maximum or inflection point in the evaluation profile precedes the divergence of outcomes. Notably, in all games, this divergence occurs before strategic tension reaches its absolute peak. This ordering suggests that the game’s outcome is largely determined during the ascent toward peak tension, before the subsequent reduction in tension differentiates decisive games from draws. In chess terms, foundational advantages—such as positional pressure, space, or superior piece coordination—are established as the board becomes increasingly complex.
@@
Attack-defense balance: When distinguishing between attack and defense links, we find that the surplus of attack links over defense links, or “attack-defense balance” as shown in Fig. 4, moves inversely with the tension profile. It initially decreases during the early, more defense-oriented phase of the game. As the game progresses, attack links gradually outnumber defense links. The curve is shifted towards more positive values in human games compared to AI games.
@@
Figure 5: Tension loads for different Elo ratings and Stockfish depths. (a) We represent tension loads (sum of tension values up to ply 150) for rapid games (turquoise dots), blitz games (green dots), classical games played by top human players (blue star), and games between Stockfish and Leela from official TCEC tournaments (red star), all associated with estimated Elo ratings. Each dot/star represents the average across 1,200 games. The blue and red stars were derived from the data analyzed in Fig. 2. For reference, blitz games are played with very short time controls (typically 3–5 minutes per player), rapid games with intermediate time controls (around 10–25 minutes per player), and classical games with long time controls (often 60 minutes or more per player) allowing deeper strategic play. (b) We represent tension loads from games between versions of Stockfish at fixed depth levels (orange dots), computed from a smaller sample of 120 games per dot. The two panels also show the $R^{2}$ values of the corresponding linear regressions.
While engines can maintain higher tension for a longer duration compared to humans, does this difference depend on skill level? We consider the cumulative tension up to ply 150 (Tension load) across different Elo ratings for human players in Fig. 5a  [^3] In principle, Stockfish has a search depth parameter that serves as a proxy for skill level. Leela has a time limit parameter during which it is permitted to assess the board state and make a decision. At a few dozen search depths, the tension load is limited to about 450, similarly for Leela at a few seconds. Since the games that are shown go far beyond these limited parameter values with supercomputers, this suggests that algorithmic complexity increases the ability to handle tension.. Tension changes substantially with Elo, showing an approximately linear dependence in human games. Moreover, across all Elo values, rapid games consistently exhibit slightly higher tension loads than blitz games. This is consistent with the idea that longer time controls allow human players to create and manage more complex positions [5]. The effect is even more pronounced in classical games, where players have significantly more time to make decisions, resulting in higher tension and reflecting the chess community’s concern that time-constrained players experience larger performance drops on complex decisions and therefore tend to choose safer, less strategically risky continuations [5, 7]. In contrast, elite classical play is often characterized by keeping simplification at bay and maintaining a complex struggle, where grandmasters sustain tense, fluid positions that test the opponent’s stamina and depth of calculation over many hours [22, 30, 12].
@@ III Discussion
AC led the definition of the tension metric, with contributions from the other authors, and carried out the code development, numerical analyses, and figure generation. EDL performed simulations on a computing cluster using Stockfish at high depths. VDPS provided the original inspiration for the project, supervised its development at all stages, and independently cross-checked the results. All authors participated in the discussions and contributed to writing the manuscript.
@@ Acknowledgements
We thank V. Latora for valuable discussions, particularly regarding the inclusion of Fig. 4, and M. Barthelemy for useful discussions and for sharing his code, which we used as a baseline for our analysis. This work was partly funded by the City of Vienna, Municipal Department 7, and the Federal Ministry of the Republic of Austria for Climate Action, Environment, Energy, Mobility, Innovation, and Technology as part of the project GZ 2023-0.841.266. It was also partially funded by the Austrian Science Fund (FWF) 10.55776/ESP127.